# Do Comments and Expertise Still Matter? An Experiment on Programmers’ Adoption of AI-Generated JavaScript Code

### v1

@@ 4.2 Hypotheses Testing
We employed linear regression to examine the influence of ($A_{1}$) presence of comments and ($B_{1}$) development expertise on (Y) latent code similarity, using a data analysis approach consistent with previous studies [23, 47]. We hypothesize that the level of development expertise exerts a moderation effect [48] on the influence of comments on programmers’ adoption of AI-generated code because more experienced programmers may rely less on comments to understand and integrate the code, while less experienced programmers may find comments crucial for comprehension and adoption. To explore the potential moderation effect, we added an interaction term into the linear regression models. The interaction term was calculated as the product of ($A_{1}$) presence of comments and ($B_{1}$) development expertise, denoted as $A_{1}*B_{1}$.
We firstly included ($A_{1}$) presence of comments, ($B_{1}$) development expertise and the interaction term $A_{1}*B_{1}$ into the linear regression. The overall model significantly predicted ($Y$) latent code similarity, $F(3,173)=2.724,p=0.46$. The coefficient table for the factors $A_{1}$, $B_{1}$, and $A_{1}*B_{1}$ is shown in Table 7. The results demonstrated that ($A_{1}$) the presence of comments could significantly increase programmers’ adoption of AI-generated code ($\beta=0.336,p=0.027$), supporting hypothesis $H_{1}$. We anecdotally observed that comments were copied in the majority of cases and that participants also tended to copy function signatures. A systematic analysis of these behaviors is part of our future work.
As showcased in Table 7, neither ($B_{1}$) development expertise nor the interaction term between ($B_{1}$) development expertise and ($A_{1}$) presence of comments had a significant influence on code adoption ($p>0.05$). This result indicated that the level of development expertise did not moderate the influence of ($A_{1}$) presence of comments. Programmers tended to adopt more AI-generated code when comments were present, regardless of their development expertise. The insignificance of $B_{1}$ and $A_{1}*B_{1}$ suggested that ($B_{1}$) development expertise did not diminish the positive effect of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code. Therefore, we reject hypothesis $H_{2}$.
@@
Variable Coefficient Significance
($A_{1}$) presence of comments 0.336 0.027
($B_{1}$) development expertise -0.113 0.253
($A_{1}*B_{1}$) interaction term -0.015 0.922
To further confirm and validate the results in Table 7, we added all control variables, including ($B_{2}$) gender, ($B_{3}$) age, ($B_{4f}$) fear of AI code generators, and ($B_{4a}$) attitudinal acceptance of AI code generators, into our initial linear regression model. The coefficients and significance of all variables included in the second model are presented in Table 8. Table 8 indicated that ($A_{1}$) presence of comments was a significant predictor of (Y) latent code similarity, while ($B_{1}$) development expertise and the interaction term $A_{1}*B_{1}$ remained insignificant. The second linear regression model confirmed the results of the initial regression model: even when accounting for all control variables, the presence of comments increases programmers’ adoption of AI-generated code regardless of programmers’ development expertise. Table 8 also showed that all control variables were insignificant, including participants’ attitudes towards AI code generators. The discrepancy between self-reported attitudes and actual behavior demonstrates the limitation of overly relying on self-reported intention of adopting AI coding tools, further highlighting the importance of adopting objective measurements of programmers’ code adoption behavior.
@@
Variable Coefficient Significance
($A_{1}$) presence of comments 0.354 0.022
($B_{1}$) development expertise -0.138 0.176
($A_{1}*B_{1}$) interaction term -0.021 0.889
($B_{2}$) gender 0.117 0.490
($B_{3}$) age 0.009 0.401
($B_{4f}$) fear of AI code generators 0.003 0.921
($B_{4a}$) attitudinal acceptance of AI code generators -0.047 0.280
@@ 5 Discussion and Implications
Our research has shown that comments, an essential element in facilitating code comprehension, continue to play a significant role in influencing programmers’ adoption of AI-generated code. The presence of comments significantly encourages programmers to adopt AI-generated code, regardless of their development expertise. Both novice and expert programmers tended to adopt more AI-generated code when it included comments.
For novice programmers, the rationale for the influence of comments could be explained in the following way: the presence of comments enhances code comprehensibility, and better code comprehensibility results in a stronger intention to adopt AI-generated code.
@@
The comprehensibility of AI-generated code plays an important role in participants’ adoption of AI-generated code. Previous studies have shown that many users perceived code understandability as a limitation of AI coding tools [5, 49, 10]. GitHub Copilot user studies discovered that users found it difficult to understand the code generated by Copilot [5, 49], and they did not use Copilot to perform programming tasks due to the difficulty in understanding the Copilot-generated code [5]. Liang et al. conducted a large-scale survey to evaluate the factors that influence the adoption and rejection of AI coding assistants [10]. Their study indicated that 16% of the participants rated ”I don’t understand the code written by code generation tools” as ”Very important” or ”Important” when asked about reasons for not using AI coding tools [10]. The limitation of unsatisfying code comprehensibility was one of the reasons programmers did not use AI coding tools, and poorly written documentation or the absence of comments could be the reason for unsatisfying code understandability.
Our study revealed that the presence of comments can also increase the adoption of AI-generated code for expert programmers. Previous studies have discovered that expert programmers were more capable of understanding source code [23, 24]. It is reasonable to extrapolate that expert programmers were more capable of understanding the AI-generated code without comments. However, they were still more likely to adopt AI-generated code that was commented. The rationale for the influence of comments on expert programmers’ code adoption might not be fully explained by the impact of comments on code comprehensibility. One possible reason why expert programmers prefer commented code is that commented code often indicates higher code quality, and expert programmers may have more stringent requirements for code quality. Code comments could be leveraged to improve the readability, maintainability, and reliability of software [50]. Expert programmers are expected to produce high-quality code that exhibits several desirable characteristics, including enhanced readability and strong maintainability [22]. Commented code was more capable of meeting the requirements of expert programmers.
@@ 6 Limitations and Threats to Validity
Measuring AI-generated code adoption through similarity metrics can be objective but may miss some aspects of adoption. While our metrics capture the adoption of code structures and exact lines of code, they overlook the adoption of stylistic aspects, such as identifier naming styles, variable naming conventions, and overall code structuring preferences. Specifically, TP and TA ignore identifier name changes, and LP and LA assess only exact line matches. Thus, if a participant adopts naming styles of the AI-generated code, like the camel case naming style, or follows function decomposition patterns present in the AI-generated code, our metrics will not capture this. Additionally, our metrics also fail to capture the situation where programmers might unconsciously adopt AI-generated code structures, for instance, having a preference for arrow functions instead of traditional function expressions when working with JavaScript. Future research could incorporate style similarity analysis and investigate whether prolonged exposure to AI-generated code affects programmers’ personal coding styles over time.
@@ 7 Conclusion and Contributions
Secondly, we proposed novel objective measurements for programmers’ adoption of AI-generated code, which are valuable complements and alternatives to traditional self-reported data. Rather than directly inquiring of participants about the degree to which they adopted AI-generated code, we employed a more objective approach. We measured the degree of code adoption by calculating the similarity between the participants’ submitted code and the AI-generated code. We evaluated code similarity from two distinct dimensions: the entity perspective and the code similarity level. This new approach could provide more objective and comprehensive insights about programmers’ adoption of AI coding tools. We have also provided the replication package [^2] The data preprocessing scripts, raw data, and the data analysis results can be found at: https://github.com/ArthurLCW/CodeResponseAnalysisScript to help future researchers utilize our measurements of code adoption.
@@
Thirdly, we developed a custom survey application that facilitated coding tasks [^3] To access the survey app, please visit: https://survey.changwen-software-engineering.xyz; Source code of the survey app can be found at: https://github.com/ArthurLCW/CodeResponseSurveyWebsite. This survey application incorporated a code editor with advanced features, including syntax highlight and auto-completion, providing a coding experience similar to using VSCode [27]. This allowed participants to complete coding tasks in a familiar environment. Future survey-based research can adapt and extend this application to suit their specific requirements, especially if the surveys include coding components.

### v2

@@ 4.2 Hypotheses Testing
We employed linear regression to examine the influence of ($A_{1}$) presence of comments and ($B_{1}$) development expertise on (Y) latent code similarity, using a data analysis approach consistent with previous studies [23, 49]. We hypothesize that the level of development expertise exerts a moderation effect [50] on the influence of comments on programmers’ adoption of AI-generated code because more experienced programmers may rely less on comments to understand and integrate the code, while less experienced programmers may find comments crucial for comprehension and adoption. To explore the potential moderation effect, we added an interaction term into the linear regression models. The interaction term was calculated as the product of ($A_{1}$) presence of comments and ($B_{1}$) development expertise, denoted as $A_{1}\times B_{1}$.
We firstly included ($A_{1}$) presence of comments, ($B_{1}$) development expertise and the interaction term $A_{1}\times B_{1}$ into the linear regression. The overall model significantly predicted ($Y$) latent code similarity, $F(3,173)=2.724,p=0.046$. The coefficient table for the factors $A_{1}$, $B_{1}$, and $A_{1}\times B_{1}$ is shown in Table 6. The results demonstrated that ($A_{1}$) the presence of comments could significantly increase programmers’ adoption of AI-generated code ($\beta=0.336,p=0.027$), supporting hypothesis $H_{1}$ [^2] To address task-specific variability and uneven dropouts across the five programming tasks, we re-analyzed the data using a mixed-effects model, which included a random intercept for task as the random effect and retained the presence of comments as the fixed effect. The re-analysis confirmed our core findings: the presence of comments remained a significant predictor of AI code adoption ($p=0.019$), supporting $H_{1}$. We anecdotally observed that comments were copied in the majority of cases and that participants also tended to copy function signatures. A systematic analysis of these behaviors is part of our future work.
As showcased in Table 6, neither ($B_{1}$) development expertise nor the interaction term between ($B_{1}$) development expertise and ($A_{1}$) presence of comments had a significant influence on code adoption ($p>0.05$). This result indicated that the level of development expertise did not moderate the influence of ($A_{1}$) presence of comments [^3] To further validate robustness, we re-estimated the model using years of programming experience (highest-loading indicator of development expertise) instead of the composite variable $B_{1}$. Results confirmed stability: presence of comments remained significant ($p=0.037$); other variables remained insignificant (programming experience: $p=0.801$; interaction term: $p=0.561$).. Programmers tended to adopt more AI-generated code when comments were present, regardless of their development expertise. The insignificance of $B_{1}$ and $A_{1}\times B_{1}$ suggested that ($B_{1}$) development expertise did not diminish the positive effect of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code. Therefore, we reject hypothesis $H_{2}$.
@@
Variable Coefficient Standardized Coefficient Significance
($A_{1}$) presence of comments 0.336 0.168 0.027
($B_{1}$) development expertise -0.113 -0.113 0.253
($A_{1}\times B_{1}$) interaction term -0.015 -0.010 0.922
To further confirm and validate the results in Table 6, we added all control variables, including ($B_{2}$) gender, ($B_{3}$) age, ($B_{4f}$) fear of AI code generators, and ($B_{4a}$) attitudinal acceptance of AI code generators, into our initial linear regression model. The coefficients and significance of all variables included in the second model are presented in Table 7. Table 7 indicated that ($A_{1}$) presence of comments was a significant predictor of (Y) latent code similarity, while ($B_{1}$) development expertise and the interaction term $A_{1}\times B_{1}$ remained insignificant. The second linear regression model confirmed the results of the initial regression model: even when accounting for all control variables, the presence of comments increases programmers’ adoption of AI-generated code regardless of programmers’ development expertise. Table 7 also showed that all control variables were insignificant, including participants’ attitudes towards AI code generators. The discrepancy between self-reported attitudes and actual behavior demonstrates the limitation of overly relying on self-reported intention of adopting AI coding tools, further highlighting the importance of adopting objective measurements of programmers’ code adoption behavior.
@@
Variable Coefficient Standardized Coefficient Significance
($A_{1}$) presence of comments 0.354 0.177 0.022
($B_{1}$) development expertise -0.138 -0.138 0.176
($A_{1}\times B_{1}$) interaction term -0.021 -0.014 0.889
($B_{2}$) gender 0.117 0.008 0.490
($B_{3}$) age 0.009 -0.084 0.401
($B_{4f}$) fear of AI code generators 0.003 0.054 0.921
($B_{4a}$) attitudinal acceptance of AI code generators -0.047 0.067 0.280
4.3 Post-hoc Analysis
While the latent code similarity variable captures the overall effects of comments and development expertise, we conducted post-hoc analyses to examine their more nuanced effects on individual code similarity metrics—specifically, line-level and token-sequence-level similarities. We re-ran the linear regression models with TA, TP, LA, and LP as dependent variables separately. The results of these regression models are presented in C.
These analyses show that for line-level similarities (LA and LP), the presence of comments ($A_{1}$) is not a significant predictor, whereas development expertise ($B_{1}$) is. In contrast, for token-sequence-level similarities (TA and TP), comment presence ($A_{1}$) is significant, while development expertise ($B_{1}$) is not. This suggests that the global effects observed in our primary model were primarily driven by token-sequence-level similarities.
As introduced in subsection 3.4.3, line-level code similarities primarily reflect direct copy-paste behavior, whereas token-sequence-level similarities capture the adoption of structural patterns. A possible explanation for the differing predictors is that less experienced programmers may be more inclined to directly copy and paste lines from the AI-generated solution due to limited confidence or an inability to devise their own approach. In contrast, more experienced programmers may be able to construct a rough solution independently and thus adopt the structural logic of the AI-generated code rather than its exact lines. This behavioral difference may explain why development expertise predicts line-level similarity but not token-level similarity.
These findings underscore the importance of being cautious in the operationalization of “code similarity,” as different aspects of similarity may be predicted by different factors. Nevertheless, given our study’s focus on the broader effects of comments and expertise, we continue to interpret and discuss the latent code similarity variable in the main manuscript, as it captures multiple dimensions of code similarity.
@@ 5 Discussion and Implications
Our research demonstrates that comments—an essential element in facilitating code comprehension—remain significant in shaping programmers’ adoption of AI-generated code within the context of lightweight JavaScript function coding. The presence of comments significantly encourages programmers to adopt AI-generated code, regardless of their development expertise. Both novice and expert programmers tended to adopt more AI-generated code when it included comments.
For novice programmers, the influence of comments on the adoption of AI-generated code can be understood through the Unified Theory of Acceptance and Use of Technology (UTAUT) [14]. According to UTAUT, technologies that are perceived as easier to use are more likely to be adopted [14]. Commented code is typically easier to comprehend, and this enhanced comprehensibility can strengthen users’ intention to adopt AI-generated solutions.
@@
The comprehensibility of AI-generated code plays an important role in participants’ adoption decisions. Prior studies have shown that many users perceive code understandability as a limitation of AI coding tools [5, 51, 10]. GitHub Copilot user studies discovered that users found it difficult to understand the code generated by Copilot [5, 51], which led some to refrain from using it for programming tasks [5]. In a large-scale survey, Liang et al. found that 16% of participants rated ”I don’t understand the code written by code generation tools” as ”Very important” or ”Important” when asked about reasons for not using AI coding tools [10]. These findings align with UTAUT’s emphasis on effort expectancy—users are less likely to adopt a technology they perceive as difficult to use [14]. When AI-generated code lacks comprehensibility, it increases cognitive load and reduces perceived ease of use, thereby discouraging adoption.
Our study revealed that the presence of comments can also increase the adoption of AI-generated code for expert programmers. Previous studies have discovered that expert programmers were more capable of understanding source code [23, 24]. It is reasonable to extrapolate that expert programmers were more capable of understanding the AI-generated code without comments. However, they were still more likely to adopt AI-generated code that was commented. This indicates that the role of comments extends beyond mere comprehensibility. For expert programmers, the presence of comments may serve as a signal of higher code quality—an aspect they prioritize more strongly due to their elevated expectations and professional standards. Code comments could be leveraged to improve the readability, maintainability, and reliability of software [52]. Expert programmers are expected to produce high-quality code that exhibits several desirable characteristics, including enhanced readability and good maintainability [22]. Commented code was more capable of meeting the requirements of expert programmers. In this case, their adoption may be more driven by performance expectancy in UTAUT [14], specifically the belief that using well-documented AI-generated code will enhance their capability to execute high-quality, maintainable development work.
@@ 6 Limitations and Threats to Validity
Measuring AI-generated code adoption through similarity metrics can be objective but may miss some aspects of adoption. While our metrics capture the adoption of code structures and exact lines of code, they overlook the adoption of stylistic aspects, such as identifier naming styles and function declaration styles. Specifically, TP and TA ignore identifier name changes, and LP and LA assess only exact line matches. Thus, if a participant adopts naming styles of the AI-generated code, like the camel case naming style, our metrics will not capture this. Additionally, our metrics also overlook cases where programmers may adopt the function declaration style used in AI-generated code, such as favoring arrow functions over traditional function expressions when working with JavaScript.
Another potential constraint on using code similarity as a measure of code adoption is the risk of chance overlap—defined as coincidental similarities between the AI-generated solution and a participant’s code that emerge not from deliberate adoption of the AI output, but because the participant independently developed a solution resembling the AI’s. While our analysis of 20 randomly selected code pairs (comparing solutions before and after exposure to the AI) suggested that chance overlap generally does not undermine the metric’s validity, residual possibilities remain: in rare cases, such unintended coincidences could potentially compromise the metric’s accuracy in reflecting actual adoption.
@@ 7 Conclusion and Future Work
Secondly, we proposed novel behavioral measurements for programmers’ adoption of AI-generated code, which are valuable complements and alternatives to traditional self-reported data. Rather than directly inquiring of participants about the degree to which they adopted AI-generated code, we employed a behavioral approach. We measured the degree of code adoption by calculating the similarity between the participants’ submitted code and the AI-generated code. We evaluated code similarity from two distinct dimensions: the entity perspective and the code similarity level. This new approach could provide behavioral and comprehensive insights about programmers’ adoption of AI coding tools. We have also provided the replication package [^4] The data preprocessing scripts, raw data, and the data analysis results can be found at: https://github.com/ArthurLCW/CodeAdoptionResearch/tree/main/analysis to help future researchers utilize our measurements of code adoption.
@@
Thirdly, we developed a custom survey application that facilitated coding tasks [^5] To access the survey app, please visit: https://survey.changwen-software-engineering.xyz; Source code of the survey app can be found at: https://github.com/ArthurLCW/CodeAdoptionResearch/tree/main/survey-web. This survey application incorporated a code editor with advanced features, including syntax highlight and auto-completion, providing a coding experience similar to using VSCode [27]. This allowed participants to complete coding tasks in a familiar environment. Future survey-based research can adapt and extend this application to suit their specific requirements, especially if the surveys include coding components.
Future studies could build on our work by addressing its limitations and further validating our findings. One potential direction is adapting our survey application to support additional programming languages. Our app currently integrates the Monaco Editor—the same editor used in VSCode—which provides advanced features like syntax highlighting natively for JavaScript and TypeScript. To offer similar support for other languages, such as Python and Swift, future researchers would need to integrate the Language Server Protocol (LSP, a standardized interface that enables code editors to communicate with language-specific servers, providing consistent advanced features like auto-completion and syntax checking [^6] More information regarding LSP: https://microsoft.github.io/language-server-protocol/) with the Monaco Editor.
Additionally, future studies could expand our survey application to better reflect real-life programming scenarios. In our current study, participants are provided with pre-generated AI solutions, based on the assumption that they would typically query AI only once for a LeetCode problem. However, in real-world settings, developers often interact with AI tools iteratively, generating multiple solutions using different prompts. To more accurately simulate these interactions, future researchers should enhance the survey app to support on-demand generation of AI solutions at runtime, with the option to include or exclude comments depending on the experimental design.
Future researchers could extend our survey application to collect and analyze richer behavioral data, such as the number of copy and paste operations, to gain deeper insights into how programmers adopt AI-generated code. Specifically, the application could be enhanced to log user interactions, including copy actions, paste actions, and selection changes, to better capture participants’ behavior during code adoption.
In addition to further validating our study results, future research could deepen our understanding of how comments influence programmers’ adoption of AI-generated code. While there is ample prior research on the impact of comments on code comprehension [18, 19, 20, 21, 52], future studies could build on these designs to investigate how varying types, lengths, and qualities of comments affect the extent to which programmers adopt AI-generated code.
Appendix A PCA Results for Development Expertise
We examined the linearity assumption [48] through the correlation matrix in Table 8. As presented in Table 8, all observed variables have at least one correlation with another observed variable greater than 0.3. The correlation coefficient is widely used as an indicator of linearity, where the correlation coefficient larger than 0.3 indicates linearity [55]. The matrix table indicated that every variable had a linear relationship with some other variables.
Table 8: Correlation matrix of observed developer factors
$B_{11\text{sub}}$ $B_{11\text{obj}}$ $B_{12\text{sub}}$ $B_{12\text{obj}}$ $B_{13\text{sub}}$ $B_{13\text{obj}}$
$B_{11\text{sub}}$ 1.000 0.571 0.445 0.470 0.547 0.527
$B_{11\text{obj}}$ 0.571 1.000 0.288 0.748 0.372 0.337
$B_{12\text{sub}}$ 0.445 0.288 1.000 0.503 0.347 0.323
$B_{12\text{obj}}$ 0.470 0.748 0.503 1.000 0.318 0.290
$B_{13\text{sub}}$ 0.547 0.372 0.347 0.318 1.000 0.724
$B_{13\text{obj}}$ 0.527 0.337 0.323 0.290 0.724 1.000
We tested the sampling adequacy assumption using the Kaiser-Meyer-Olkin (KMO) measure and Bartlett’s test of sphericity. The overall KMO value was 0.703, exceeding the 0.5 threshold [46], with all individual KMO values also above 0.5 (see Table 9), suggesting sufficient shared variance for factor analysis. Bartlett’s test of sphericity was significant ($p<0.001$) [47], indicating that the correlation matrix was not an identity matrix. Together, these results confirmed the data’s suitability for factor analysis.
Table 9: KMO measures of observed developer factors
Observed Variable KMO Measure
$B_{11\text{sub}}$ 0.824
$B_{11\text{obj}}$ 0.632
$B_{12\text{sub}}$ 0.669
$B_{12\text{obj}}$ 0.636
$B_{13\text{sub}}$ 0.742
$B_{13\text{obj}}$ 0.734
PCA revealed two components with eigenvalues greater than 1, explaining 54.729% and 18.169% of the total variance, respectively. Based on the scree plot criterion [48, 56], we retained only the first component, which appears before the inflection point in the scree plot (Fig. 9). This decision was supported by the high explained variance (54.729%) and the interpretability of the component. We applied Varimax rotation with Kaiser normalization [57, 58], and the rotation matrix (Table 10) showed that the first component adequately represented all observed variables. We then reran PCA to extract a single component and recalculated the loadings (Table 11). The resulting component aligned with the concept of Development Expertise ($B_{1}$), as it positively loaded on general programming experience, language proficiency, and domain-specific knowledge.
Figure 9: Scree plot of principal components for observed programmer factors
Table 10: Rotation matrix with component loadings
Observed Variable Component 1 Component 2
$B_{11\text{sub}}$ 0.572 0.315
$B_{11\text{obj}}$ 0.922 0.103
$B_{12\text{sub}}$ 0.158 0.898
$B_{12\text{obj}}$ 0.853 0.189
$B_{13\text{sub}}$ 0.205 0.886
$B_{13\text{obj}}$ 0.564 0.589
Table 11: Component matrix for development expertise
Observed Variable Coefficient
$B_{11\text{sub}}$ 0.751
$B_{11\text{obj}}$ 0.813
$B_{12\text{sub}}$ 0.748
$B_{12\text{obj}}$ 0.758
$B_{13\text{sub}}$ 0.722
$B_{13\text{obj}}$ 0.635
Table 12 reports the communalities for each observed variable included in the PCA. These values represent the proportion of variance in each variable accounted for by the extracted principal component [59]. Most variables show high communalities (e.g., $B_{12\text{obj}}$ = 0.861; $B_{13\text{sub}}$ = 0.827), indicating they are well explained by the component. This supports the validity of summarizing these variables into a single latent factor representing development expertise ($B_{1}$).
Table 12: Communalities for observed variables in development expertise
Observed Variable Communality
$B_{11\text{sub}}$ 0.665
$B_{11\text{obj}}$ 0.763
$B_{12\text{sub}}$ 0.427
$B_{12\text{obj}}$ 0.861
$B_{13\text{sub}}$ 0.827
$B_{13\text{obj}}$ 0.832
Appendix B PCA Results for Code Similarity
Similar to the process for generating development expertise ($B_{1}$), we first tested the linearity and sampling adequacy assumptions. Linearity was assessed using the correlation matrix in Table 13. All code similarity metrics showed at least one correlation above 0.6, with several exceeding 0.9 (e.g., between TP and TA; LP and LA), indicating strong inter-correlations [55]. These high correlations satisfy the linearity assumption and suggest the metrics are closely related. The strong inter-correlations may be attributed to the short length of the code solutions (typically under 25 lines), which limits structural variability and amplifies similarity across metrics.
Table 13: Correlation matrix of code similarity metrics
TP TA LP LA
TP 1.000 0.962 0.669 0.628
TA 0.962 1.000 0.687 0.688
LP 0.669 0.687 1.000 0.969
LA 0.628 0.688 0.969 1.000
After testing linearity, we assessed sampling adequacy. The overall Kaiser-Meyer-Olkin (KMO) measure was 0.527, with all individual KMO values above 0.5 (Table 14), indicating sufficient shared variance for dimensionality reduction [46]. Bartlett’s test of sphericity was significant ($p<0.001$), suggesting the correlation matrix was not an identity matrix [47]. Together, these results confirmed that the sampling adequacy assumption was met.
Table 14: KMO measures of code similarity metrics
Code Similarity Metric KMO Measure
TP 0.515
TA 0.538
LP 0.537
LA 0.516
PCA revealed one principal component with an eigenvalue above 1, explaining 82.550% of the total variance. The scree plot (Fig. 10) confirmed that only the first component should be retained, as the second point marked the inflection point. This component captured the core variance across all four code similarity metrics and was interpretable as a unified latent variable. We used it as the dependent variable, with its component loadings shown in Table 15.
Figure 10: Scree plot of principal components for code similarity metrics
Table 15: Component matrix for latent code similarity
Code Similarity Metric Coefficient
TP 0.915
TA 0.905
LP 0.896
LA 0.918
We examined the communalities to assess how much of each code similarity metric’s variance is captured by the extracted principal component. As shown in Table 16, all four code similarity metrics had high extraction values, ranging from 0.803 to 0.843. This indicates that the retained component effectively represents the shared variance across these metrics, supporting the dimensionality reduction.
Table 16: Communalities for code similarity metrics
Code Similarity Metric Extraction
TP 0.838
TA 0.818
LP 0.803
LA 0.843
Appendix C Post-Hoc Regression Analyses on Individual Code Similarity Metrics
We conducted post-hoc regression analyses to examine how development expertise ($B_{1}$) and the presence of comments ($A_{1}$) influence individual code similarity metrics. Specifically, we analyzed: token-sequence-level similarity from the participant’s perspective (TP), token-sequence-level similarity from the AI’s perspective (TA), line-level similarity from the participant’s perspective (LP), and line-level similarity from the AI’s perspective (LA).
We employed the same linear regression model described in Subsection 4.2, replacing the latent code similarity variable with each of the four individual similarity metrics as dependent variables. The regression results are summarized in Tables 17 through 20.
Table 17: Regression Results with TA as Dependent Variable
Variable Coefficient Standardized Coefficient $p$-value
($A_{1}$) Presence of comments 0.157 0.208 0.006
($B_{1}$) Development expertise -0.012 0.012 0.835
($A_{1}\times B_{1}$) Interaction term -0.021 0.005 0.900
Table 18: Regression Results with TP as Dependent Variable
Variable Coefficient Standardized Coefficient $p$-value
($A_{1}$) Presence of comments 0.118 0.156 0.041
($B_{1}$) Development expertise -0.002 -0.004 0.964
($A_{1}\times B_{1}$) Interaction term -0.001 -0.003 0.980
Table 19: Regression Results with LA as Dependent Variable
Variable Coefficient Standardized Coefficient $p$-value
($A_{1}$) Presence of comments 0.093 0.138 0.064
($B_{1}$) Development expertise -0.070 -0.207 0.034
($A_{1}\times B_{1}$) Interaction term -0.011 -0.020 0.833
Table 20: Regression Results with LP as Dependent Variable
Variable Coefficient Standardized Coefficient $p$-value
($A_{1}$) Presence of comments 0.070 0.109 0.147
($B_{1}$) Development expertise -0.068 0.004 0.032
($A_{1}\times B_{1}$) Interaction term -0.211 0.009 0.930